OpenStudy (anonymous):

I need the answers to these questions quickly, please: Choose the equation of the line passing through the point (-4, -2)and parallel to y = 1/4 x + 2. y = 1/4 x - 1 y = 4x - 1 y = 1/4 x - 3 y = 4x - 3 Choose the equation of the line passing through the point (-1, 7)and perpendicular to y = -1/4x - 5. y = 4x + 3 y = -4x - 11 y = -4x - 3 y = 4x + 11

OpenStudy (anonymous):

y= 1/4 x for the first one y=4x +11 for the second one

OpenStudy (anonymous):

y = 1/4 x? That isn't one of the answers on it

OpenStudy (anonymous):

y=1/4 x -1 Since it has the same slope and it passes thru the given point.

OpenStudy (anonymous):

Okay, thanks ^_^

OpenStudy (ash2326):

@eliassaab please explain how did you get the answer

OpenStudy (anonymous):

(y-(-2))/(x-(-4)) = 1/4 (y+2)/(x+4) = 1/4 4(y+2) = x+4 4y+8 = x+4 4y = x-4 y = 1/4x-1 parallel lines have the same slope. For perpendicular lines, the product of the slopes is -1, so the slope of the line perpendicular to the 2nd one is 4 For both of these, you use the slope formula to get the equation.

OpenStudy (anonymous):

For the second one the product of the slopes must be -1 so the answer is either choice one or choice 4 to pass thru the point, one has to choose 4